OpenStudy (anonymous):

Which table best represents the steps to solve the equation - x over 6 - 3 = -9? @TheRealMeeeee @katalina13

OpenStudy (anonymous):

Step 1: - x over 6 = -9 - 3 The same number -3 can be added to both sides of the equation Step 2: - x over 6 = -12 Adding -9 and -3 on the right hand side Step 3: x = -12 • -6 The same number -6 can be multiplied on both sides of the equation Step 4: x = 72 Multiplying -12 and -6 on the right hand side Step Justification Step 1: - x over 6 = -9 + 3 The same number +3 can be added to both sides of the equation Step 2: - x over 6 = -6 Adding -9 and 3 on the right hand side Step 3: x = -6 -6 The same number -6 can be added on both sides of the equation Step 4: x = -12 Adding -6 and -6 on the right hand side Step Justification Step 1: - x over 6 = -9 + 3 The same number + 3 can be added to both sides of the equation Step 2: - x over 6 = -6 Adding -9 and 3 on the right hand side Step 3: x = -6 ÷ -6 The same number -6 can be divided on both sides of the equation Step 4: x = 1 Dividing -6 by -6 on the right hand side Step Justification Step 1: - x over 6 = -9 + 3 The same number + 3 can be added to both sides of the equation Step 2: - x over 6 = -6 Adding -9 and 3 on the right hand side Step 3: x = -6 • -6 The same number -6 can be multiplied on both sides of the equation Step 4: x = 36 Multiplying -6 and -6 on the right hand side

OpenStudy (therealmeeeee):

dang

OpenStudy (therealmeeeee):

i cant help u on this one lmao

OpenStudy (anonymous):

ur not trying lol

OpenStudy (anonymous):

@bibby

OpenStudy (therealmeeeee):

LMAO no sirry bob

OpenStudy (anonymous):

what??

OpenStudy (therealmeeeee):

grrr i cant ddo this i dont feel reading it

OpenStudy (anonymous):

@tester97

OpenStudy (anonymous):

yd u give me another medal?

OpenStudy (therealmeeeee):

yes u neeed some

OpenStudy (therealmeeeee):

watch wat im bout to do

OpenStudy (anonymous):

what

OpenStudy (anonymous):

omg stop

OpenStudy (anonymous):

haha thats cheating